OpenStudy (anonymous):

Which of the following shows the correct order of the numbers below? (1/2)squared,2/3 , 0.45 1)(1/2)squared < 0.45 < 2/3 2)(1/2)squared < 2/3< 0.45 3)0.45 < (1/2)squared < 2/3

OpenStudy (anonymous):

thank you please forgive me

OpenStudy (anonymous):

tell @peachpi to help me and that i am sorry

OpenStudy (rushwr):

\[\left( \frac{ 1 }{ 2 } \right)^{2} = \frac{ 1 }{ 4 } = 0.25 \] Did u get this part ?

OpenStudy (anonymous):

it is asking for the order

OpenStudy (anonymous):

i hate math

OpenStudy (rushwr):

okai so u got that part right?

OpenStudy (rushwr):

so 2/3 = 0.66 right?

OpenStudy (anonymous):

oh right yea yea

OpenStudy (anonymous):

i am so tired this morning

OpenStudy (rushwr):

now we have all of the numbers in decimals. SO arrange these number: 0.25, 0.45 and 0.66

OpenStudy (rushwr):

U got it right? What's ur answer ?See maths is fun and easy?

OpenStudy (anonymous):

0.66?,0.45?,0.25?

OpenStudy (anonymous):

oh never mind i found it out

OpenStudy (rushwr):

What's ur answer so ?

OpenStudy (anonymous):

A

OpenStudy (rushwr):

Exactly !!!!

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

my internet is slow where i live (Sardis) and plus it is raining so i will close out most of m windows and get to work before i get kicked off the internet

OpenStudy (anonymous):

i mean my windows

OpenStudy (rushwr):

oki Medal me may be ?